OpenStudy (anonymous):

solve the system using any algebraic method 0.25x+0.5y=7.5 0.4x+0.5y=9 Show work for a medal :)

OpenStudy (dumbcow):

Use elimination Multiply everything by 100, that will get rid of the decimals :) Multiply top equation by -1 Add equations to eliminate y -25x -50y = -750 40x +50y = 900 ---------------- 15x + 0 = 150 15x = 150 x = 10

OpenStudy (anonymous):

so whats y

OpenStudy (anonymous):

take the top equation and put it in terms of y: .5y = 7.5 - .25x y = 15 - .5x plug into second equation and solve for x 0.4x + 0.5(15 - .5x) = 9 0.4x + 7.5 - .25x = 9 0.15x = 1.5 x = 10 plug x into any equation to find y: .5y = 7.5 - .25(10) y = 15 - 2.5 = 12.5

OpenStudy (dumbcow):

i figured you could figure that out substitute x=10 into one of original equations

OpenStudy (dumbcow):

y is not 12.5 what do you think?

OpenStudy (anonymous):

x=10 and y=10

OpenStudy (dumbcow):

alright 40x + 50y = 900 x = 10 400 + 50y = 900 50y = 500 y = 10

OpenStudy (anonymous):

y = 15 - .5(10) y = 15 - 5 y = 10